OpenStudy (anonymous):

Find the period, range, phase shift and vertical shift of f(x)=9/2csc(8x+4pi/7)+6

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Here, on the first page on the bottom they show you how to analyze it into the parameters you're interested in =) https://www.cedarvalleycollege.edu/FutureStudents/StudentServices/TutoringServices/FAQandResources/Math/Pictures/Transformations%20sec%20and%20csc.pdf

OpenStudy (anonymous):

I'm a little confused by that

OpenStudy (anonymous):

wait would the period be 2pi/8?

OpenStudy (anonymous):

Ok, let's see. they say that if we have $$ A \cdot \csc(B \cdot x + C) + D $$Then: \(|A|\) is our amplitude \(\frac{2\pi}{B}\) is our period. \(-\frac{C}{B}\) is our phase shift \(D\) is our vertical shift and finally \(-A\) is our X axis flip, which we're not asked here

OpenStudy (anonymous):

So yes, the period would be \(\frac{2\pi}{8}\)

OpenStudy (anonymous):

what about finding range?

OpenStudy (anonymous):

look at the first page on the graph above. you can see that they say Range: \( y \le -1\) or \(y \ge 1\) for csc(x)

OpenStudy (anonymous):

if c is 4pi/7 and be is 8 then what would be the phase shift?

OpenStudy (anonymous):

b*

OpenStudy (anonymous):

Well, \(-\frac{C}{B} = - \frac{ \frac{4\pi}{7}}{8} = -\frac{4\pi}{7 \cdot 8} = -\frac{\pi}{14}\)